# FuzzySQL: Uncovering Hidden Vulnerabilities in DBMS Special Features with LLM-Driven Fuzzing

### v1

@@ Abstract.
Traditional database fuzzing techniques primarily focus on syntactic correctness and general SQL structures, leaving critical yet obscure DBMS features, such as system-level modes (e.g., GTID), programmatic constructs (e.g., PROCEDURE), advanced process commands (e.g., KILL), largely underexplored. Although rarely triggered by typical inputs, these features can lead to severe crashes or security issues when executed under edge-case conditions. In this paper, we present FuzzySQL, a novel LLM-powered adaptive fuzzing framework designed to uncover subtle vulnerabilities in DBMS special features. FuzzySQL combines grammar-guided SQL generation with logic-shifting progressive mutation, a novel technique that explores alternative control paths by negating conditions and restructuring execution logic, synthesizing structurally and semantically diverse test cases. To further ensure deeper execution coverage of the back end, FuzzySQL employs a hybrid error repair pipeline that unifies rule-based patching with LLM-driven semantic repair, enabling automatic correction of syntactic and context-sensitive failures. We evaluate FuzzySQL across multiple DBMSs, including MySQL, MariaDB, SQLite, PostgreSQL and Clickhouse, uncovering 37 vulnerabilities, 7 of which are tied to under-tested DBMS special features. As of this writing, 29 cases have been confirmed with 9 assigned CVE identifiers, 14 already fixed by vendors, and additional vulnerabilities scheduled to be patched in upcoming releases. Our results highlight the limitations of conventional fuzzers in semantic feature coverage and demonstrate the potential of LLM-based fuzzing to discover deeply hidden bugs in complex database systems.
Database security; Fuzzing; AI for security; Language validity
@@ 1. Introduction
To address these limitations, we present FuzzySQL, a novel adaptive fuzzing framework driven by LLMs and designed to systematically uncover vulnerabilities in the under-tested feature space of modern DBMSs. FuzzySQL introduces the following key innova-
1CREATE TABLE t0 (xxx);
2ALTER DATABASE test_db READ ONLY = 1;
3ANALYZE TABLE t0 UPDATE HISTOGRAM ON v1;
4RESET BINARY LOGS AND GTIDS;
5INSERT INTO t0 VALUES (xxx);
@@
tions. First, FuzzySQL targets feature-specific SQL constructs by combining grammar-guided template expansion with LLM-powered instantiation (Section 3.1), enabling systematic coverage of vendor-specific and version-dependent DBMS features that are difficult to model using fixed rules. Second, to overcome the challenges of semantic modeling, FuzzySQL performs semantic exploration through a unified design that integrates progressive mutation with automated error repair (Sections 3.2 and 3.3). Specifically, it prioritizes semantic diversity by recombining logic fragments across seed programs and applying conditional logic shifts, while automatically repairing high-value test cases that fail due to missing objects or unmet semantic preconditions. This design enables FuzzySQL to construct semantically valid and diverse SQL sequences that reach deep, feature-dependent execution paths. Finally, FuzzySQL maintains the global execution context throughout fuzzing and performs replay-based analysis to identify and isolate compound failures that arise only under specific internal DBMS states (Section 3.4). This paper makes the following key contributions:
@@
  • Discovering real-world vulnerabilities across multiple DBMSs. We evaluate FuzzySQL on MySQL, MariaDB, SQLite, and ClickHouse, uncovering 37 vulnerabilities, of which 29 have been confirmed by developers and 7 are associated with under-tested DBMS features. These results highlight FuzzySQL’s ability to expose semantic and feature-specific vulnerabilities consistently missed by existing approaches.
@@ 2.1. Under-Tested Features in Modern DBMSs
While traditional database fuzzers primarily target general-purpose SQL statements such as SELECT, INSERT, and CREATE TABLE, modern DBMSs expose a broader attack surface through system-level features that manipulate internal state and metadata. Listing 1 presents a minimal SQL sequence that triggers a persistent crash in MySQL by misusing such special-purpose statements.
These features include operations for toggling system modes (e.g., READ ONLY), managing replication state (e.g., BINARY LOGS and GTID), and maintaining internal statistics (e.g., UPDATE HISTOGRAM). Although invoked less frequently than standard DML and DDL statements, they directly interact with core DBMS subsystems and global execution state. Each statement in Listing 1 is syntactically valid and semantically reasonable in isolation. However, their specific combination exposes a subtle control-path vulnerability in MySQL’s metadata lock management. In particular, toggling READ ONLY mode and updating table histograms interact with metadata locks maintained by the dictionary subsystem, while resetting binary logs alters session-level replication state. When followed by a regular INSERT, these interactions violate internal lock invariants and trigger an assertion failure in the metadata lock checker.
More critically, triggering this vulnerability leads to persistent state corruption that survives across subsequent executions. Even after restarting the server, later queries result in immediate crashes, indicating a lasting impact on internal state rather than a transient fault. This example demonstrates that under-tested DBMS features can interact in unexpected ways, producing severe and state-dependent failures that are difficult to expose through traditional statement-level fuzzing.
@@ 2.3. LLM: Opportunities and Challenges
LLMs enable semantic understanding of complex SQL constructs, allowing them to generate coherent multi-statement SQL sequences, infer implicit dependencies, and synthesize missing objects (e.g., creating a PROCEDURE when a CALL is issued). This reduces reliance on manually crafted rules in DBMS fuzzing and supports more comprehensive testing of system-specific features. However, LLMs also introduce unique challenges in the fuzzing context. Due to their exposure to well-formed data during pretraining and instruction tuning, they exhibit a strong bias toward generating syntactically correct, semantically plausible, and conventionally safe inputs. Specific design choices are required to guide LLMs toward semantic reasoning over schema context, rather than defaulting to generic low-risk outputs.
A notable characteristic of LLMs is hallucination, where models generate syntactically or semantically invalid content relative to the underlying system or schema. While typically undesirable, hallucinations can be beneficial for fuzzing, as they help explore unexpected behaviors and may expose corner-case bugs. In the fuzzing context, hallucination acts as a feature that promotes exploration beyond normative inputs. We further illustrate this phenomenon in Section 5.4.
@@ 2.4. Goals of FuzzySQL
FuzzySQL aims to shift the paradigm of database fuzzing from traditional syntax-driven or random mutation strategies toward a semantically guided approach that emphasizes deeper program logic and broader exploration of system features. We provide an anonymized artifact repository to support reproducibility, as described in the Open Science Appendix C.
@@ 3. Methodology
To uncover deep and stateful bugs in modern DBMSs, FuzzySQL adopts a unified fuzzing pipeline that integrates structured generation, semantic mutation, automated repair, and replay-based validation, as illustrated in Figure 1. Rather than generating isolated SQL statements, FuzzySQL continuously evolves multi-statement programs under a persistent execution context, enabling exploration of feature-specific and state-dependent behaviors. Fuzzing begins with grammar-guided SQL template expansion, followed by LLM-based instantiation under the current schema state to produce executable and semantically meaningful inputs. Executed test cases that yield new coverage are retained and further evolved via progressive mutation, which recombines statement sequences and applies logic-level transformations to explore alternative control paths. When execution fails due to syntactic or semantic inconsistencies, FuzzySQL applies a multi-stage repair pipeline to salvage high-value inputs instead of discarding them. Finally, detected crashes are validated through replay-guided execution and minimized to concise, reproducible PoCs that capture state-dependent failures.
@@ 3.1.1. Grammar Template Expansion
1alterTable -> ALTER ONLINE TABLE [tableName] [alterSpecification] [partitionDefinitions]
2alterTable -> ALTER TABLE [tableName] [alter-Specification],[alterSpecification]
@@ 3.3.1. Error Taxonomy
To ensure the soundness of our repair strategy, we conduct a quantitative analysis of error messages returned by MySQL, a representative and widely deployed DBMS. From 12 hours of grammar-guided SQL generation, we collect 65,591 errors spanning 184 distinct types (out of 1,960 documented MySQL client error codes). Syntax-related errors account for 29.61% of all messages, while the top 50 semantic error types contribute 68.44%. These high-frequency semantic errors are manually analyzed and mapped to dedicated rule-based or intelligent semantic-aware repair strategies, while the remaining long-tail categories are handled via LLM-based semantic repair.
@@ 3.3.2. Syntax-Aware Filtering
Before invoking any repair strategy, FuzzySQL performs error tagging to map DBMS error messages back to the corresponding SQL statements. This is achieved through log analysis and keyword alignment, allowing precise localization of faulty fragments for subsequent filtering or repair.
For syntax-related errors, FuzzySQL applies structural filtering to prioritize repair-worthy candidates. Specifically, SELECT statements with abstract syntax tree (AST) depth fewer than three are discarded as overly simplistic and unlikely to exercise deep execution logic. In contrast, statements involving DBMS-specific features (e.g., GTID, PROCEDURE, INSTALL COMPONENT) are always retained, as these often reside in under-tested paths and are more likely to expose non-trivial bugs. In addition, FuzzySQL filters out errors which are caused by duplicate object definitions or unsupported features rather than input errors. These include errors like Table 'XX'already exists, or This version of [DBMS] does not yet support 'XXX'. Filtering these cases early allows FuzzySQL to focus computational resources on repairing inputs with genuine structural or semantic promise.
@@
To efficiently handle frequent and structurally shallow error cases, FuzzySQL incorporates a rule-based repair layer that operates without requiring deep semantic modeling. This module targets predictable error patterns that can be corrected through localized adjustments—such as replacing invalid literals, reformatting malformed clauses, or adjusting configuration parameters—based solely on pattern matching and heuristic rewriting. Compared to semantic reasoning approaches, rule-based repair is lightweight and DBMS-agnostic. For example, errors such as ’Cannot get geometry object from data you send to the GEOMETRY field’ can be resolved by adjusting the literal assignment to conform to the expected data type. It avoids the overhead of parsing syntax trees or inferring dependencies, making it portable across dialects. More importantly, it enables fast recovery from common failures without invoking LLMs, thereby conserving LLM query budget for more complex, high-value repairs. By pre-staging repairable cases into this rule-based stage, FuzzySQL ensures high-throughput processing of grammar-generated inputs.
@@ 3.3.4. Intelligent Semantic-Aware Repair
For complex test failures rooted in semantic inconsistencies, such as undefined objects, unmet preconditions, or illegal feature combinations—FuzzySQL invokes a LLM to perform intelligent and context-sensitive repair. Unlike deterministic fixes, this stage leverages LLMs’ ability to interpret program intent and synthesize structurally valid SQL fragments to recover test cases that would otherwise be discarded.
The repair process begins by grouping remaining errors by type (e.g., invalid object reference, missing prerequisite, constraint violation), each associated with its corresponding faulty SQL fragment and execution context. For each case, FuzzySQL constructs a structured repair prompt that includes the SQL snippet, relevant schema history, the DBMS error message, and a natural language instruction describing the repair objective. Based on this prompt, the LLM generates a patch, either by inserting missing definitions (e.g., creating tables or functions) or rewriting invalid clauses. These transformations are injected directly into the test case for re-evaluation. For instance, when encountering an error such as PROCEDURE xxx does not exist, FuzzySQL constructs a prompt that instructs the LLM to synthesize a valid CREATE PROCEDURE template, placing it before the failing invocation. This effectively guides the repair process toward restoring missing database objects and leads to the generation of an interesting PROCEDURE.
In particular, LLM repairs are highly effective at recovering DBMS-specific constructs and operation sequences. This error-repair loop forms an iterative Chain-of-Thought (COT), where the model refines test cases based on error messages and repair strategies. This final repair phase enables FuzzySQL to recover high-value test inputs that exhibit deep structural complexity, significantly increasing the coverage of under-tested DBMS behaviors.

### v2

@@ Abstract.
Traditional database fuzzing techniques primarily focus on syntactic correctness and general SQL structures, leaving critical yet obscure DBMS features, such as system-level modes (e.g., GTID), programmatic constructs (e.g., PROCEDURE), advanced process commands (e.g., KILL), largely underexplored. Although rarely triggered by typical inputs, these features can lead to severe crashes or security issues when executed under edge-case conditions. In this paper, we present FuzzySQL, a novel LLM-powered adaptive fuzzing framework designed to uncover subtle vulnerabilities in DBMS special features. FuzzySQL combines grammar-guided SQL generation with logic-shifting progressive mutation, a novel technique that explores alternative control paths by negating conditions and restructuring execution logic, synthesizing structurally and semantically diverse test cases. To further ensure deeper execution coverage of the back end, FuzzySQL employs a hybrid error repair pipeline that unifies rule-based patching with LLM-driven semantic repair, enabling automatic correction of syntactic and context-sensitive failures. We evaluate FuzzySQL across multiple DBMSs, including MySQL, MariaDB, SQLite, PostgreSQL and Clickhouse, uncovering 64 vulnerabilities, 27 of which are tied to under-tested DBMS special features. As of this writing, 60 cases have been confirmed with 9 assigned CVE identifiers, 31 already fixed by vendors, and additional vulnerabilities scheduled to be patched in upcoming releases. Our results highlight the limitations of conventional fuzzers in semantic feature coverage and demonstrate the potential of LLM-based fuzzing to discover deeply hidden bugs in complex database systems.
Database security, Fuzzing, AI for security, Language validity
@@ 1. Introduction
1 CREATE TABLE t0 (xxx);
2 ALTER DATABASE test_db READ ONLY = 1;
3 ANALYZE TABLE t0 UPDATE HISTOGRAM ON v1;
4 RESET BINARY LOGS AND GTIDS;
5 INSERT INTO t0 VALUES (xxx);
@@
To address these limitations, we present FuzzySQL, a novel adaptive fuzzing framework driven by LLMs and designed to systematically uncover vulnerabilities in the under-tested feature space of modern DBMSs. FuzzySQL introduces the following key innovations. First, FuzzySQL targets feature-specific SQL constructs by combining grammar-guided template expansion with LLM-powered instantiation (Section 3.1), enabling systematic coverage of vendor-specific and version-dependent DBMS features that are difficult to model using fixed rules. Second, to overcome the challenges of semantic modeling, FuzzySQL performs semantic exploration through a unified design that integrates progressive mutation with automated error repair (Sections 3.2 and 3.3). Specifically, it prioritizes semantic diversity by recombining logic fragments across seed programs and applying conditional logic shifts, while automatically repairing high-value test cases that fail due to missing objects or unmet semantic preconditions. This design enables FuzzySQL to construct semantically valid and diverse SQL sequences that reach deep, feature-dependent execution paths. Finally, FuzzySQL maintains the global execution context throughout fuzzing and performs replay-based analysis to identify and isolate compound failures that arise only under specific internal DBMS states (Section 3.4). This paper makes the following key contributions:
@@
  • Discovering real-world vulnerabilities across multiple DBMSs. We evaluate FuzzySQL on MySQL, MariaDB, SQLite, and ClickHouse, uncovering 64 vulnerabilities, of which 60 have been confirmed by developers and 27 are associated with under-tested DBMS features. These results highlight FuzzySQL’s ability to expose semantic and feature-specific vulnerabilities consistently missed by existing approaches.
@@ 2.1. Under-Tested Features in Modern DBMSs
Traditional database fuzzers mainly focus on general-purpose SQL statements such as SELECT, INSERT, and CREATE TABLE. However, modern DBMSs expose a much broader attack surface through SQL-accessible system features that manipulate internal metadata, runtime modes, and persistent execution state. Listing 1 shows a minimal SQL sequence that triggers a persistent crash in MySQL by exercising such under-tested feature interactions.
In this paper, we use under-tested DBMS features to refer to functionalities that are vendor-specific, version-evolving, or operationally uncommon, yet directly affect core DBMS subsystems. Representative categories include configuration and state toggles (e.g., READ ONLY), replication and logging controls (e.g., BINARY LOGS and GTID), and optimizer or statistics maintenance operations (e.g., UPDATE HISTOGRAM). Thus, the challenge is broader than covering rare syntax alone: these features typically require non-trivial semantic preconditions and often interact with sensitive internal subsystems such as metadata management, replication logic, and statistics maintenance.
The example in Listing 1 illustrates this challenge. Each statement is syntactically valid and semantically reasonable in isolation, yet their particular combination exposes a subtle vulnerability in MySQL’s metadata lock management. Specifically, toggling READ ONLY mode and updating table histograms affect metadata locks maintained by the dictionary subsystem, while resetting binary logs modifies session-level replication state. When followed by a regular INSERT, these accumulated effects violate internal lock invariants and trigger an assertion failure in the metadata lock checker. More importantly, the bug is state-dependent and persistent. Once triggered, it corrupts internal state in a way that survives subsequent executions: even after restarting the server, later queries may still crash immediately. This example shows that feature-related DBMS bugs are hard to expose not simply because such statements are less frequently generated, but because they require the fuzzer to jointly cover feature-specific constructs, satisfy their semantic preconditions, and preserve the execution history in which harmful interactions emerge.
@@ 2.3. LLM: Opportunities and Challenges
LLMs provide an effective complement to traditional DBMS fuzzing because they can reason about SQL at the semantic level rather than only at the syntactic level. This capability is particularly useful for exercising under-tested DBMS features, whose correct invocation often depends on implicit object dependencies, valid statement ordering, and feature-specific execution context. In practice, LLMs can generate coherent multi-statement SQL sequences, infer missing prerequisites, and synthesize auxiliary objects or definitions (e.g., creating a PROCEDURE before issuing a CALL). By doing so, they reduce the need for manually crafted feature-specific rules and improve the ability of the fuzzer to reach semantically constrained execution paths.
However, using LLMs for fuzzing also introduces challenges. Since these models are trained and aligned primarily on well-formed examples, they are naturally inclined to produce syntactically correct, semantically plausible, and conventionally safe outputs. This tendency is useful for standard code generation, but less ideal for fuzzing, where valuable test cases often arise from unusual feature combinations, edge-case dependencies, or execution contexts that depart from typical usage. As a result, LLMs must be guided carefully so that they reason over schema context and execution history instead of defaulting to generic low-risk completions.
A related issue is hallucination, where the model generates content that does not fully match the actual DBMS dialect, schema, or runtime context. Although such deviations are usually undesirable, they can occasionally introduce exploratory diversity beyond normative SQL patterns and thereby help expose unexpected feature interactions or corner-case failures. For this reason, FuzzySQL does not treat LLM output as inherently trustworthy or inherently harmful; instead, it uses downstream execution feedback and repair to retain useful exploratory behaviors while filtering out low-value failures. We further illustrate this phenomenon through case studies in Section 5.4.
@@ 2.4. Goals of FuzzySQL
Overall, FuzzySQL aims to shift DBMS fuzzing beyond purely syntax-driven generation or random mutation toward a semantically guided workflow that explores both deeper program logic and broader system functionality. More importantly, FuzzySQL is not merely a general-purpose LLM-assisted SQL fuzzer. Its design explicitly addresses three obstacles that disproportionately hinder the discovery of feature-related DBMS bugs: (1) insufficient coverage of vendor- and version-specific feature constructs; (2) unmet semantic preconditions that prevent such features from being exercised effectively; and (3) hidden state interactions that span multiple statements or executions. Accordingly, FuzzySQL combines feature-oriented grammar expansion and LLM instantiation, semantics-aware mutation and repair, and replay-guided validation under persistent execution context.
@@ 3. Methodology
To uncover deep and stateful bugs in modern DBMSs, FuzzySQL adopts a unified fuzzing pipeline that integrates structured generation, semantic mutation, automated repair, and replay-based validation, as illustrated in Figure 1. Rather than generating isolated SQL statements, FuzzySQL continuously evolves multi-statement programs under a persistent execution context, enabling exploration of feature-specific and state-dependent behaviors. Fuzzing begins with grammar-guided SQL template expansion, followed by LLM-based instantiation under the current schema state to produce executable and semantically meaningful inputs. Executed test cases that yield new coverage are retained and further evolved via progressive mutation, which recombines statement sequences and applies logic-level transformations to explore alternative control paths. When execution fails due to syntactic or semantic inconsistencies, FuzzySQL applies a multi-stage repair pipeline to salvage high-value inputs instead of discarding them. Finally, detected crashes are validated through replay-guided execution and minimized to concise, reproducible PoCs that capture state-dependent failures. Each stage is designed to address a distinct obstacle in uncovering feature-related DBMS bugs: grammar-guided generation expands the reachable feature space, mutation and repair satisfy non-trivial semantic preconditions, and replay-guided validation exposes failures whose triggers are distributed across persistent execution history.
@@ 3.1.1. Grammar Template Expansion
1 alterTable -> ALTER ONLINE TABLE [tableName] [alterSpecification] [partitionDefinitions]
2 alterTable -> ALTER TABLE [tableName] [alter-Specification],[alterSpecification]
@@ 3.3.1. Error Taxonomy
To ensure the soundness of our repair strategy, we conduct a quantitative analysis of error messages returned by MySQL, a representative and widely deployed DBMS. From 12 hours of grammar-guided SQL generation, we collect 65,591 errors spanning 184 distinct types (out of 1,960 documented MySQL client error codes). Syntax-related errors account for 29.61% of all messages, while the top 50 semantic error types contribute 68.44%. These high-frequency semantic errors are manually analyzed and mapped to dedicated rule-based or intelligent semantic-aware repair strategies.
@@ 3.3.2. Syntax-Aware Filtering
Before invoking any repair strategy, FuzzySQL first classifies execution errors according to the taxonomy introduced in Section 3.3.1. This step determines whether an error should be repaired or filtered out directly. In particular, FuzzySQL focuses repair efforts on syntax and semantic errors that still indicate structurally or semantically promising SQL inputs, while excluding failures that are unlikely to contribute useful execution behaviors.
To this end, FuzzySQL maintains a lightweight, DBMS-specific set of error patterns. Each pattern is defined by an error code, a regular-expression-style keyword matcher over the DBMS error message, and an associated action. When an error matches a filtering rule, FuzzySQL labels it as SAF and discards the corresponding input from further repair. Typical examples include duplicate-definition errors (e.g., messages containing already exists) and unsupported-feature errors (e.g., Method ... is not supported). These failures usually indicate environmental conflicts or feature absence rather than repairable input defects, and thus provide little value for subsequent mutation or semantic correction. For the remaining syntax-related failures, FuzzySQL applies structural filtering to prioritize repair-worthy candidates. Specifically, statements with fewer than twenty tokens are more likely to have an abstract syntax tree (AST) depth of fewer than three; such statements are discarded as overly shallow and unlikely to exercise deep execution logic.
This filtering strategy allows FuzzySQL to avoid wasting repair budget on low-value failures while preserving inputs that are more likely to trigger meaningful feature interactions or deeper internal behaviors. In practice, the goal of this stage is not compiler-grade diagnostic precision, but to narrow the repair scope sufficiently so that subsequent repair modules can focus on high-value candidates.
@@
After syntax- and error-aware filtering, FuzzySQL applies a rule-based repair (RBR) layer to handle error cases that admit predictable, pattern-driven fixes. Similar to the filtering stage, this module is triggered by DBMS-specific error patterns defined through error codes, message keywords, and associated repair actions. Once an error is mapped to RBR, FuzzySQL invokes the corresponding repair operator to transform the faulty SQL block before re-execution.
The key idea of RBR is to correct recurring failures through lightweight rewriting, without invoking full semantic reasoning. Depending on the matched error type, the repair may insert auxiliary statements, adjust the local statement context, or attach a recommended SQL template for subsequent repair stages. For example, when MySQL reports an error indicating that a table was not locked with LOCK TABLES, FuzzySQL invokes an operator that prepends UNLOCK TABLES; to the faulty SQL block. Likewise, when MariaDB reports that a referenced table does not exist, FuzzySQL can invoke a createtable operator that inserts a recommended CREATE TABLE template near the failing statement, thereby supplying structured context for later semantic repair.
Compared with directly invoking the LLM on every failing input, RBR provides a lightweight intermediate layer for high-frequency, structurally regular errors. Its role is not limited to fully repairing the input in one step; in many cases, it also prepares a more informative and constrained context for the subsequent intelligent semantic-aware repair stage. This design allows FuzzySQL to recover quickly from common failures while reserving LLM-based reasoning for cases that genuinely require deeper semantic inference.
@@ 3.3.4. Intelligent Semantic-Aware Repair
After syntax- and error-aware filtering and rule-based repair, FuzzySQL forwards the remaining repairable failures to an intelligent semantic-aware repair stage powered by an LLM. This stage handles two categories of cases: (1) inputs that remain unresolved after earlier lightweight repair, and (2) failures that are not filtered out as low-value errors but do not match any available rule-based repair pattern. Instead of discarding such inputs prematurely, FuzzySQL attempts to recover them through context-sensitive semantic repair, so that structurally rich and potentially high-value test cases can continue to participate in fuzzing.
The rationale for this stage is that many DBMS failures, especially those involving under-tested features, cannot be corrected through local rewriting alone. Such failures often arise from incomplete semantic dependencies, including missing objects, unmet feature preconditions, incompatible session state, or invalid cross-statement ordering. These dependencies are difficult to enumerate exhaustively as hand-written rules, particularly across heterogeneous DBMS dialects and evolving feature sets. Therefore, FuzzySQL employs an LLM as a practical mechanism for inferring missing execution context and restoring semantically valid feature invocations from the failing SQL fragment and DBMS diagnostic message.
For each unresolved case, FuzzySQL constructs a structured repair prompt from the current SQL block, relevant schema or execution context, and the associated DBMS error message. The model is then asked to revise the test case so that the target feature or behavior can be exercised under a semantically valid context. Depending on the failure mode, the generated patch may introduce missing definitions, restore prerequisite statements, or rewrite incompatible clauses while preserving the original testing intent as much as possible. The repaired result is then re-injected into the fuzzing loop for re-execution.
Importantly, this repair process is inherently iterative. A single repair attempt may only partially resolve the original issue, and subsequent execution or mutation may further change the semantic context in which the test case is evaluated. As a result, FuzzySQL does not treat semantic repair as a one-shot correction step, but as a feedback-driven refinement loop in which newly observed DBMS diagnostics continuously inform the next repair decision. Conceptually, this design is related to chain-of-thought-style decomposition: instead of requiring the model to recover the entire valid execution context in a single step, FuzzySQL incrementally guides repair through successive error signals and evolving execution history. Notably, this iterative behavior arises at two levels: within the repair stage itself, where one repair may expose another unresolved dependency, and across the broader fuzzing loop, where subsequent mutations may alter the execution context and create new semantic requirements.
This stage is designed not merely to improve input validity, but to preserve difficult test cases that are more likely to exercise deeper control paths, feature interactions, and state-sensitive behaviors. In this sense, intelligent semantic-aware repair serves as the final recovery layer of FuzzySQL  it prevents semantically incomplete but high-value inputs from being discarded too early, thereby substantially improving the exploration of under-tested DBMS features.